OpenStudy (anonymous):

A car travels 100 meters in 5 seconds. What is the speed of the car? speed = distance over time; 1 mile = 1609 meters 8.9 miles per hour 16 miles per hour 20 miles per hour 44.7 miles per hour

OpenStudy (haseeb96):

1 mile = 1609 meters Speed = distance /time speed = 100 /5 = 20 m/s now 44.7 miles per hour 44.7 miles multiply with 1609 =71922 meters 1 hour = 3600 secs 71922/3600=19.98 =20 m/s D option is correct